OpenStudy (anonymous):

Given the equation y − 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in standard form. −3/4x + y = 10 3x − 4y = −40 y =3/4 x + 12 y =3/4 x + 10

OpenStudy (anonymous):

\[ y − 4 =3/4 (x + 8) =3/4 x + 6\\ \] Can you finish it?

OpenStudy (anonymous):

\[ y=\frac 3 4 x +10 \]

OpenStudy (anonymous):

The last choice.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw